Explain how you would determine the degree of polymerization using only NMR.
Answer #1

The degree of polymerization (DP) or number of repeating units of a polymer is defined as the average number of
structural units per polymer chain.. From 1H NMR spectrum,by comparing the proton peak intensity of a known moiety (typically an end-group(s) with a known number of protons) to that of the repeating chain unit of interest can give DP. An equation is derived for calculating the value of DP, which is as follows.

We know that intensity or area of a proton signal in 1H NMR is proportional to amount of that species present in sample (number of molecules of that species times its molecular weight.) From here, we have

a_{i} = k \times n_{i}\times m_{i}

Rearranging for k,

k = \frac{a_{i}}{n_{i}\times m_{i}}

Here, for a species i, ai is the area or intensity of the 1H NMR peak; ni is the number of repeating units; mi is the number of protons; and k is the constant, because number of molecules of that species is proportional to number of repeating units and molecular weight of that species is proportional to its number of protons.

Consider the 1H NMR signals of two different moieties x and y,

k_{x}= \frac{a_{x}}{n_{x}\times m_{x}}

k_{y}= \frac{a_{y}}{n_{y}\times m_{y}}

But, in a given polymer kx = ky , so

\frac{a_{x}}{n_{x}\times m_{x}} = \frac{a_{y}}{n_{y}\times m_{y}}

where ax is the area or intensity of the 1H NMR peak of moiety x; nx is the number of repeating units of moiety x; mx is the number of protons of moiety x; ay is the area or intensity of the 1H NMR peak of moiety y; ny is the number of repeating units of moiety y; and my is the number of protons of moiety y.

Rearranging  nx,

n_{x} = \frac{a_{x}\times m_{y}\times n_{y} }{a_{y}\times m_{x}}

where nx can be used to assess the polymer’s DP or number of repeating units.

Consequently, Mn can be calculated by substituting for n

M_{n} =n\times M_{o} + M_{e}

where n is the number of repeating units or DP; Mo is the molecular weight of one repeating unit, and Me is the combined molecular weight of the end-groups.

Example :

Consider the spectrum of MPEG-NH2. The protons of the methoxy (CH3O?) and amino (?NH2) end-groups resonated as a singlet and broad triplet, respectively, at 3.25 and 2.71 ppm (J = 5.6 Hz), whereas the prominent peak of the oxyethylene protons (?(OCH2CH2)n?) was centered at 3.52 ppm.

Determination of the Number of Repeating Units

To calculate the number of repeating units (n) in the MPEG-NH2 chain, the peak areas of CH3O– (OMe, ? 3.25) and ?(OCH2CH2)n– (OEt, ? 3.52) were obtained from the 1H NMR spectrum

n_{MPEG} = \frac{a_{OEt}\times m_{OMe}\times n_{OMe} }{a_{OMe}\times m_{OEt}}

= (133.1*3*1)/(0.9*4)

= 110.92 \approx 111.